What are the number of gram atoms in 63.5 g of copper


1
Expert's answer
2021-08-22T07:29:58-0400

n = m / M

M (Cu) = 63.5460 g/mol

n (Cu) = 63.5/63.5 = 1 mol

1 mol of matter contains 6.022×1023 elementary entities of the substance.

63.5 g copper = 6.023×1023 atoms of copper.